# A Semantic Invariant Robust Watermark for Large Language Models

### v1

@@ 1 Introduction
Text watermarking techniques for LLMs usually embed specific information during text generation to allow high-accuracy detection of LLM-generated text. The mainstream approach for embedding such information is to add extra watermark logits on top of the logits generated by the LLM. For example, Kirchenbauer et al. (2023) divide the vocabulary into red and green lists and increase the scores for the green tokens as the watermark logits. However, current watermarking algorithms cannot possess both attack robustness (robustness to modifications of the watermarked text) and security robustness (robustness against watermark cracking). For example, Zhao et al. (2023) demonstrates that global fixed watermark logits enhance attack robustness, yet they compromise security due to vulnerability in word frequency analysis (Sadasivan et al., 2023). Essentially, in current watermark algorithms Kirchenbauer et al. (2023); Zhao et al. (2023); Wang et al. (2023), the watermark logits for each token depend on its preceding tokens. As the number of required preceding tokens increases, watermarking complexity rises, leading to reduced attack robustness but increased security robustness.
@@
To resolve the aforementioned trade-off, we propose a semantic invariant watermarking algorithm that achieves reasonable attack and security robustness. The core motivation is generating watermark logits for each token based on the preceding tokens’ semantics rather than their token IDs. Thus, semantically invariant text modifications do not alter the watermark logits, while the diversity of text semantics increases watermark complexity and guarantees security against watermark cracking. Specifically, to extract semantically invariant features, we use an auxiliary LLM encoder (e.g., BERT) to extract semantic embeddings of the preceding tokens. We then train a small watermark network to transform these embeddings into corresponding watermark logits. The training objective of the watermark network is to ensure a high correlation between the similarities of the output watermark logits and input text embeddings. Also, it’s imperative that the watermark logits exhibit sufficient diversity and are unbiased for each token. To achieve these goals, two training objectives are adopted: a similarity loss and a normalization loss. For the similarity loss, to ensure the diversity of the watermark logits, we first rescale the similarity values between text embeddings to range from -1 to 1, and then make the similarity of the generated watermark logits fit this similarity. To ensure unbiased token selection and achieve bimodal scores of the watermark logits, our normalization loss centers the mean of each row and column within a batch of watermark logits to zero, while making the absolute value of each entry as close as possible. We detect watermarks by averaging the watermark logits across all tokens. If the average is significantly above zero, a watermark is present.
@@
In summary, the contributions of this work are as follows:
  • We propose the first semantically invariant robust watermarking algorithm, which effectively detects watermarks under various semantically invariant perturbations.
  • Our algorithm successfully navigates the trade-off between attack robustness and security robustness that plagued previous methods, achieving high performance in both dimensions.
  • We innovatively propose a watermark model that adeptly transforms semantic embeddings into watermark logits.
@@ 2 Related work
Text watermarking technology embeds imperceptible patterns within textual content and employs specialized detection techniques to ascertain the presence of a watermark (Abdelnabi & Fritz, 2021; Yang et al., 2022). Currently, there are two approaches to watermarking text: post-processing after text generation, and incorporating watermarks during the text generation of LLM.
@@
Post-processing methods make minor semantic-preserving alterations to the text, most commonly lexical substitution-based approaches. For instance, Qiang et al. (2023) embedded watermark information by employing a paraphraser-based lexical substitution, ensuring the text’s original semantics remained intact. Yoo et al. (2023) began by selecting stable words to position the watermark and then used a masked language model to generate candidate words, thereby facilitating the embedding and extraction of watermarks. Munyer & Zhong (2023) generated candidate texts through word2vec-based synonym replacements, subsequently comparing semantic similarities among candidates to finalize the watermarked text. Nonetheless, replacing individual words while only considering the unchanged meaning of those words risks degrading overall text quality. In contrast, generative modification methods like Abdelnabi & Fritz (2021) directly generate watermarked text, along with corresponding networks to recover the watermark information. However these methods can still only make minimal changes to the original text, severely limiting the freedom to insert watermarks.
@@
Compared to post-processing methods, integrating watermarks into the text generation process of current generative language models better utilizes their capabilities. Based on the stage at which watermark information is embedded, current watermarking methods for LLMs can be divided into two categories: adding watermarks during token sampling and during logits computation. Christ et al. (2023) proposed a method that embeds watermark information by presetting a random number sequence for token sampling, and detects watermarks by computing correlation between texts and the random number sequence. To improve robustness against text modifications, Kuditipudi et al. (2023) adopted Levenshtein distance for matching texts and the random number sequence, allowing watermarked texts to be detectable even after certain modifications. However, although such token sampling-based watermarking methods minimize the impact on text quality, they also greatly limit the randomness of generated texts, and only work for sampling decoders instead of other decoding methods like beam search. In contrast, another category is to add watermarks during logits computation. Kirchenbauer et al. (2023) made minor modifications to current token logits based on hashes of previous k tokens (watermark logits). Lee et al. (2023) designed a similar watermarking method for low-entropy code generation. Wang et al. (2023) enabled multi-bit watermark information via a proxy LLM. Liu et al. (2023) implemented private watermarking through parameter sharing between the watermark generator and detector. In this work, we are more concerned with watermark robustness. Zhao et al. (2023) proved that using global fixed watermark logits (k=0) in Kirchenbauer et al. (2023)’s method achieves very strong attack robustness, yet Sadasivan et al. (2023)’s work shows that lower k values make watermarks vulnerable to security robustness (easy to break). Therefore, previous watermarking methods struggle to achieve both attack and security robustness. We propose a semantically invariant robust watermarking scheme to achieve decent attack and security robustness.
@@ 3 Preliminaries
The robustness of our work encompasses two aspects: attack robustness and security robustness. Attack robustness evaluates the probability that the watermarked text can still be correctly identified after semantically invariant modifications. Security robustness evaluates the accuracy of inferring the watermark rules from the watermarked text. Once the watermarking rules are disclosed, users can easily modify the text to remove the watermark, rendering attack robustness meaningless.
@@ 4.1 Watermark Generation
As discussed in the previous sections, one of the most important steps in the watermark generation process is the generation of the watermark logits. The watermark logits for the current token are usually determined by its preceding tokens. Our goal is to construct a continuous and robust mapping from the semantics of the preceding tokens to the watermark logits. In this way, semantic-preserving modifications to the text will only cause small perturbations to the watermark logits, thus ensuring the attack robustness of the entire watermark algorithm.
To extract semantic-invariant features of the text, we utilize an embedding language model $\mathrm{E}$ (e.g. BERT). Specifically, given a sequence of preceding tokens ${\bm{t}}_{:l-1}$, we first obtain its semantic embedding ${\bm{e}}_{l}=\mathrm{E}({\bm{t}}_{:l-1})$. To transform this semantic-invariant feature to a score over the vocabulary $\mathrm{V}$, we train a specialized watermark model $\mathrm{T}$ to generate the watermark logits: $P_{\mathrm{W}}=\mathrm{T}({\bm{e}}_{l})$. The overall algorithm is described in detail in Algorithm 1. A thorough introduction to the watermark model will be provided in Section 4.2. We will also provide a detailed analysis of the algorithm’s robustness in Section 5.
@@ 4.2 Watermark Model
In this section, we introduce the watermark model in detail. The goal of the watermark model is to convert semantic embeddings into watermark logits. First, we describe several desired properties for the similarity of watermark logits: semantic-consistent broad range, unbiased token preference, and balanced score.
@@
Additionally, to further ensure the security robustness of the watermark, the watermark logits should have an unbiased token preference, meaning there should be no statistical preference for any token. This is shown in the following equation:
@@
To augment the separability of watermark logits, the output from the watermark model is further processed by the $\mathrm{tanh}$ function, yielding the final watermark logits, denoted as: $\mathrm{tanh}(k_{2}\,T(\mathbf{e}_{i}))$.
@@ 4.3 Watermark Detection
In this section, we introduce the method for detecting the watermark.
Analogous to watermark generation, the watermark logit is first computed per token via Algorithm 1. The score at the position of each token $t_{j}$ is selected. Averaging all scores gives the detection score:
$z=\frac{\sum_{j=1}^{N}P_{\mathrm{W}}^{(j)}({\bm{x}}^{prompt},\bm{t}_{:j-1})}{N}.$ (7)
Without a watermark, the expected score is 0 since the watermark logit mean is 0. When a watermark is present, the score substantially exceeds 0, making this a suitable validation approach. We utilize the z-score in the following to indicate our detection score.
@@ 6.1 Experiment settings
Dataset and Prompt: Similar to the previous works Kirchenbauer et al. (2023), we utilize the C4 dataset (Raffel et al., 2020) for data generation, taking the first 30 tokens as prompts and generating the next 200 tokens. The original C4 texts serve as human-written examples. The test objective is to distinguish between generated text and human-written text. During training of the watermark model, we utilize the WikiText-103 dataset (Merity et al., 2016) (different from C4) to generate embeddings.
Baseline and Language Model: We selected two watermarking algorithms as baselines. One is KGW-k (Kirchenbauer et al., 2023) where k is the number of preceding tokens to hash. The other is exponential minimum sampling (EXP-edit) (Kuditipudi et al., 2023), an approach that introduces watermarks through a pre-selected sequence of sampling probabilities during the sampling process and adopts edit distance-based detection to achieve strong robustness. For language models, we use LLaMA-7B (Touvron et al., 2023), OPT1.3B, and OPT2.7B (Zhang et al., 2022) for text generation. Additionally, we tested the efficacy of the watermarking algorithm under both stochastic and deterministic scenarios using sampling and beam search decoding algorithms, respectively. For embedding language models, we utilized Compositional-BERT (Chanchani & Huang, 2023) due to its superior ability to produce embeddings that better distinguish text.
@@ 6.2 Attack Robustness Analysis
In Table 1, we list the detection accuracy of our watermark algorithm and various baseline algorithms under the no-attack setting and when texts are rewritten using GPT3.5 as well as two different DIPPER settings (Krishna et al., 2023). For GPT3.5, we use the gpt-3.5-turbo-0613 version with the prompt Rewrite the following paragraph:. For DIPPER-1 the lex diversity is 60 without order diversity, and for DIPPER-2 we additionally increase the order diversity by 20. EXP-Edit cannot be used under beam search so we do not list its result.
Table 1 shows that our watermarking algorithm achieves strong robustness against all rewrite attacks. Specifically, for watermarked texts rewritten by GPT-3.5 and two DIPPER rewriters, our algorithm still achieves an average detection F1 score of 0.93. This performance is comparable to that of the KGW-1 algorithm, markedly surpassing other watermark algorithms, including KGW-2, KGW-4, and EXP-Edit. As modifications to the token in the KGW-1 algorithm only affect the label of that particular token, this represents the upper bound of attack robustness for watermark algorithms based on watermark logits. Nonetheless, subsequent experiments demonstrate that the KGW-1 algorithm possesses poor security robustness.
@@ 6.3 Security Robustness Analysis
Table 1: We compared the performance of our watermarking method with others, including KGW-k (Kirchenbauer et al., 2023) and EXP (Kuditipudi et al., 2023), using text generated by LLaMA-7B. Tests involved watermark detection accuracy under no attack, GPT3.5 rewrite attacks, and two DIPPER (Krishna et al., 2023) settings.
@@
Figure 3(a) illustrates the trade-off between attack robustness (measured by detection accuracy after rewrite) and security robustness (measured by watermark decryption accuracy) in Kirchenbauer et al. (2023)’s watermarking algorithm. A smaller k implies stronger attack robustness yet simpler watermarking rules and thus lower security robustness. Our watermarking algorithm achieves attack robustness close to KGW-1 while maintaining security robustness close to KGW-4. Although our security robustness degrades slightly with more refined domain text (L1 and L2 level classes), even for the very refined L2 class it is still far superior to KGW-3. Therefore, our watermarking algorithm successfully addresses the attack robustness versus security robustness trade-off in prior work.
@@
Figure 3: The left figure depicts the trade-off between security robustness (measured by watermark decryption accuracy) and attack robustness (measured by F1 score after rewrite) across different watermarking algorithms. The right figure shows the text quality generated by language models with different watermarking methods (measured by text perplexity).
@@ 6.4 Watermark Model Analysis
To better analyze our watermarking algorithm, Figure 2(b) shows the relationship between the similarity of text embeddings and the similarity of watermark logits generated by the watermark model. The average similarity of watermark logits is shown for each range of embedding similarity. It can be seen that the similarity of watermark logits maintains a strong correlation with the embedding similarity, especially for very similar embeddings that certainly correspond to very similar watermark logits (similarity greater than 0.8). Meanwhile, the range of watermark logit similarity is also extended to -1 to 1 (the figure only shows the average similarity in the range so -1 is not shown), which also makes the watermarking rules more diverse and robust. From the figure, the average similarity of the original sentence embeddings is 0.45, which also reflects the advantage of using Compositional-BERT, which can better distinguish dissimilar texts compared to the original BERT, helping to improve the diversity of watermarks. We will analyze in more detail in the appendix the impact of different embedding models on watermarking. Figure 2(b) also shows our normalization loss results in less distinguished watermark logits.

### v2

@@ 1 Introduction
Text watermarking techniques for LLMs usually embed specific information during text generation to allow high-accuracy detection of LLM-generated text. The mainstream approach for embedding such information is to add extra watermark logits on top of the logits generated by the LLM. For example, Kirchenbauer et al. (2023a) divide the vocabulary into red and green lists and increase the scores for the green tokens as the watermark logits. However, current watermarking algorithms cannot possess both attack robustness (robustness to modifications of the watermarked text) and security robustness, which refers to the difficulty of inferring watermarking rules from watermarked text. For example, Zhao et al. (2023) demonstrates that global fixed watermark logits enhance attack robustness, yet they compromise security due to vulnerability in word frequency analysis (Sadasivan et al., 2023). This is because the frequency of tokens from their green list is much higher compared to those in normal text, the high-frequency tokens could be simply treated as green tokens and further used to remove the watermark. Essentially, in current watermark algorithms, the watermark logits for each token depend on its preceding tokens. As the number of required preceding tokens increases, watermarking complexity rises, leading to reduced attack robustness but increased security robustness.
@@
To resolve the aforementioned trade-off, we propose a semantic invariant watermarking algorithm that achieves reasonable attack and security robustness. The core motivation is generating watermark logits for each token based on the preceding tokens’ semantics rather than their token IDs. Thus, semantically invariant text modifications do not alter the watermark logits, while the diversity of text semantics increases watermark complexity and guarantees security against watermark cracking. Specifically, to extract semantically invariant features, we use an auxiliary LLM encoder (e.g., BERT) to extract semantic embeddings of the preceding tokens. We then train a small watermark network to transform these embeddings into corresponding watermark logits. The training objective of the watermark network is to ensure a high correlation between the similarities of the output watermark logits and input text embeddings. Also, it’s imperative that the watermark logits exhibit sufficient diversity and are unbiased for each token. To achieve these goals, two training objectives are adopted: a similarity loss and a normalization loss. For the similarity loss, to ensure the diversity of the watermark logits, we first rescale the similarity values between text embeddings to range from -1 to 1, and then make the similarity of the generated watermark logits fit this similarity. To ensure unbiased token selection and achieve bimodal scores of the watermark logits, our normalization loss centers the mean of each row and column within a batch of watermark logits to zero, while making the absolute value of each entry as close as possible. During the detection phase, we compute the watermark logits for each token at its position and obtain the corresponding value. We then average the values across all tokens and determine the presence of a watermark by checking if the average is significantly greater than zero.
@@
Our contributions are as follows: (1) We propose the first semantically invariant robust watermarking algorithm, which effectively detects watermarks under various semantically invariant perturbations. (2) Our algorithm successfully navigates the trade-off between attack robustness and security robustness that plagued previous methods, achieving high performance in both dimensions. (3) We propose a watermark model that adeptly transforms semantic embeddings into watermark logits.
@@ 2 Related work
Currently, there are two approaches to watermarking text Liu et al. (2023c): post-processing after text generation, and incorporating watermarks during the text generation of LLM.
@@
Post-processing methods make minor semantic-preserving alterations to the text, most commonly lexical substitution-based approaches. For instance, Qiang et al. (2023) embedded watermark information by employing a paraphraser-based lexical substitution, ensuring the text’s original semantics remained intact. Yoo et al. (2023) began by selecting stable words to position the watermark and then used a masked language model to generate candidate words, thereby facilitating the embedding and extraction of watermarks. Munyer & Zhong (2023) generated candidate texts through word2vec-based synonym replacements, subsequently comparing semantic similarities among candidates to finalize the watermarked text. Nonetheless, replacing individual words while only considering the unchanged meaning of those words risks degrading overall text quality. In contrast, generative modification methods like Abdelnabi & Fritz (2021) directly generate watermarked text, along with corresponding networks to recover the watermark information.
@@
Based on the stage at which watermark information is embedded, current watermarking methods for LLMs can be divided into two categories: adding watermarks during token sampling and during logits computation. Christ et al. (2023) proposed a method that embeds watermark information by presetting a random number sequence for token sampling, and detects watermarks by computing correlation between texts and the random number sequence. To improve robustness against text modifications, Kuditipudi et al. (2023) adopted Levenshtein distance for matching texts and the random number sequence, allowing watermarked texts to be detectable even after certain modifications. However, although such token sampling-based watermarking methods minimize the impact on text quality, they also greatly limit the randomness of generated texts, and only work for sampling decoders instead of other decoding methods like beam search. In contrast, another category is to add watermarks during logits computation. Kirchenbauer et al. (2023a) made minor modifications to current token logits based on hashes of previous k tokens (watermark logits). Lee et al. (2023) designed a similar watermarking method for low-entropy code generation. Wang et al. (2023) enabled multi-bit watermark information via a proxy LLM. Liu et al. (2023b) implemented watermarking through parameter sharing between the watermark generator and detector. In this work, we are more concerned with watermark robustness. Zhao et al. (2023) proved that using global fixed watermark logits (k=0) in Kirchenbauer et al. (2023a)’s method achieves very strong attack robustness, yet Sadasivan et al. (2023)’s work shows that lower k values make watermarks vulnerable to security robustness (easy to break). In this work, we propose a semantically invariant robust watermarking scheme, where watermark logits are generated based on semantic information of the text, achieving decent attack and security robustness.
@@ 3 Preliminaries
The robustness of our work encompasses two aspects: attack robustness and security robustness. Attack robustness evaluates the probability that the watermarked text can still be correctly identified after semantically invariant modifications. Security robustness evaluates the accuracy of inferring the watermark rules from the watermarked text. Once the watermarking rules are disclosed, users can easily modify the text to remove the watermark, rendering attack robustness meaningless. In our work, security robustness is evaluated by employing spoofing attacks Sadasivan et al. (2023), specifically by conducting statistical analysis on watermarked texts.
@@ 4.1 Watermark Generation
As discussed in the previous sections, one of the most important steps in the watermark generation process is the generation of the watermark logits. The watermark logits for the current token are usually determined by its preceding tokens. Our goal is to construct a continuous and robust mapping from the semantics of the preceding tokens to the watermark logits. In this way, semantic-preserving modifications to the text will only cause small perturbations to the watermark logits.
To extract semantic-invariant features of the text, we utilize an embedding language model $\mathrm{E}$ (e.g. BERT). Specifically, given a sequence of preceding tokens ${\bm{t}}_{:l-1}$, we first obtain its semantic embedding ${\bm{e}}_{l}=\mathrm{E}({\bm{t}}_{:l-1})$. To transform this semantic-invariant feature to a score over the vocabulary $\mathrm{V}$, we train a specialized watermark model $\mathrm{T}$ to generate the watermark logits: $P_{\mathrm{W}}=\mathrm{T}({\bm{e}}_{l})$. The overall algorithm is described in detail in Algorithm 1. A thorough introduction to the watermark model will be provided in Section 4.2.
@@ 4.2 Watermark Model
The goal of the watermark model is to convert semantic embeddings into watermark logits. First, we describe several desired properties for the similarity of watermark logits: semantic-consistent broad range, unbiased token preference, and balanced score.
@@
Additionally, to further ensure the security robustness of the watermark, the watermark logits should have an unbiased token preference, meaning there should be no statistical preference for any token:
@@
To augment the separability of watermark logits, the output from the watermark model is further processed by the $\mathrm{tanh}$ function, yielding the final watermark logits, denoted as: $\mathrm{tanh}(k_{2}\,T(\mathbf{e}_{i}))$. After this procedure, the values of watermark logits are almost exclusively 1 or -1. This correlates with the concept of red-green tokens in the work of Kirchenbauer et al. (2023a) (denoted as KGW), indicating a fundamental similarity in principles between our approach and the KGW algorithms.
@@ 4.3 Watermark Detection
Similar to some previous works Kirchenbauer et al. (2023a), we assume the null hypothesis and compute a z-statistic. We reject the null hypothesis and detect the watermark if the average watermark logit for each token is greater than zero. The watermark logit for each token is calculated via Algorithm 1 and its average is zero and standard variation is one (as shown in Figure 2(c)), which could be represented as:
$z=\frac{\sum_{j=1}^{N}(P_{\mathrm{W}}^{(t_{j})}(x^{prompt},\bm{t}_{:j-1})-0)}{N*1}=\frac{\sum_{j=1}^{N}P_{\mathrm{W}}^{(t_{j})}(x^{prompt},\bm{t}_{:j-1})}{N}.$ (7)
Without a watermark, the expected score is 0 since the watermark logit mean is 0. When a watermark is present, the score substantially exceeds 0, making this a suitable validation approach. In our work, the values of watermark logits are almost exclusively 1 or -1, which corresponds to the concept of green-red tokens by Kirchenbauer et al. (2023a). Consequently, our detection methodology is fundamentally the same to theirs, a point we elaborate on in greater detail in the appendix D. Note that the watermarking model is optimized so that the null hypothesis is approximately true on non-watermarked text, but no guarantee can be provided (e.g. on out-of-domain text).
@@ 6.1 Experiment settings
Dataset and Prompt: Similar to the previous works Kirchenbauer et al. (2023a), we utilize the C4 dataset (Raffel et al., 2020) for data generation, taking the first 30 tokens as prompts and generating the next 200 tokens. The original C4 texts serve as human-written examples. The test objective is to distinguish between generated text and human-written text. During training of the watermark model, we utilize the WikiText-103 dataset (Merity et al., 2016) (different from C4) to generate embeddings.
Baseline and Language Model: We selected two watermarking algorithms as baselines. One is KGW-k (Kirchenbauer et al., 2023a) where k is the number of preceding tokens to hash. The other is exponential minimum sampling (EXP-edit) (Kuditipudi et al., 2023), an approach that introduces watermarks through a pre-selected sequence of sampling probabilities during the sampling process and adopts edit distance-based detection to achieve strong robustness. For language models, we use LLaMA-7B (Touvron et al., 2023), OPT1.3B, and OPT2.7B (Zhang et al., 2022) for text generation. Additionally, we tested the efficacy of the watermarking algorithm under both stochastic and deterministic scenarios using sampling and beam search decoding algorithms, respectively. For embedding language models, we utilized Compositional-BERT (Chanchani & Huang, 2023) due to its superior ability to produce embeddings that better distinguish text.
@@ 6.2 Attack Robustness Analysis
In Table 1, we list the detection accuracy of our watermark algorithm and various baseline algorithms under the no-attack setting and when texts are rewritten using GPT3.5, two different DIPPER settings (Krishna et al., 2023) and the copy-paste attack Kirchenbauer et al. (2023b). For GPT3.5, we use the gpt-3.5-turbo-0613 version with the prompt Rewrite the following paragraph:. For DIPPER-1 the lex diversity is 60 without order diversity, and for DIPPER-2 we additionally increase the order diversity by 20. For copy-paste attack, we insert 600 tokens from the origin text before the generated text.
Table 1 shows that our watermarking algorithm achieves strong robustness against all attacks. Specifically, for watermarked texts rewritten by GPT-3.5 and two DIPPER rewriters, our algorithm still achieves an average detection F1 score of 0.93. This performance is comparable to that of the KGW-1 algorithm, markedly surpassing other watermark algorithms, including KGW-2, KGW-4, and EXP-Edit. As modifications to the token in the KGW-1 only affect the label of that particular token, this represents the upper bound of attack robustness for watermark algorithms based on watermark logits. Nonetheless, subsequent experiments demonstrate that the KGW-1 algorithm possesses poor security robustness. We also achieve robustness against copy-paste attacks similar to that of the KGW series methods. Appendix G details the robustness to copy-paste and other types of attacks.
@@ 6.3 Security Robustness Analysis
Table 1: We compared the performance of our watermarking method with others, including KGW-k (Kirchenbauer et al., 2023a) and EXP (Kuditipudi et al., 2023), using text generated by LLaMA-7B. Tests involved watermark detection accuracy under no attack, GPT3.5 rewrite attacks, two DIPPER (Krishna et al., 2023) settings, the copy-paste attack as well as the emoj attack (Appendix G).
@@
Figure 3(a) illustrates the trade-off between attack robustness (measured by detection accuracy after rewrite) and security robustness (measured by watermark decryption accuracy) in Kirchenbauer et al. (2023a)’s watermarking algorithm. A smaller k implies stronger attack robustness yet simpler watermarking rules and thus lower security robustness. Our watermarking algorithm achieves attack robustness close to KGW-1 while maintaining security robustness close to KGW-4. Although our security robustness degrades slightly with more refined domain text (L1 and L2 level classes), even for the very refined L2 class it is still far superior to KGW-3. Therefore, our watermarking algorithm successfully addresses the attack robustness versus security robustness trade-off in prior work.
@@
Figure 3: The left figure depicts the trade-off between security robustness and attack robustness across different watermarking algorithms. The right figure shows the text quality generated by language models with different watermarking methods (measured by text perplexity).
@@ 6.4 Watermark Model Analysis
To better analyze our watermarking algorithm, Figure 2(b) shows the relationship between the similarity of text embeddings and the similarity of watermark logits generated by the watermark model. The average similarity of watermark logits is shown for each range of embedding similarity. It can be seen that the similarity of watermark logits maintains a strong correlation with the embedding similarity, especially for very similar embeddings that certainly correspond to very similar watermark logits (similarity greater than 0.8). Meanwhile, the range of watermark logit similarity is also extended to -1 to 1 (the figure only shows the average similarity in the range so -1 is not shown), which also makes the watermarking rules more diverse and robust. From the figure, the average similarity of the original sentence embeddings is 0.45, which also reflects the advantage of using Compositional-BERT, which can better distinguish dissimilar texts compared to the original BERT, helping to improve the diversity of watermarks. We will analyze in more detail in the appendix the impact of different embedding models on watermarking.